OpenStudy (anonymous):

MEDEAL AND FAN Solve the equation. 1. 6 6/7=n + 5 3/7 A 12 2/7 B 6 6/7 C 5 6/7 D 1 3/7 2. x-1/7=3 4/7 A 2 5/7 B 3 3/7 C 3 5/7 3. y+1/4=6 11/12 A 1 35/48 B 5 2/3 C 6 2/3 D 7 1/6 4.x-1/2=2/8 A 1/4 B 3/4 C 3 D 1 1/3

OpenStudy (anonymous):

Solve the equation by rearranging in terms of n: \[6\frac{ 6 }{ 7 }-5\frac{ 3 }{ 7 }=n\]

OpenStudy (anonymous):

turn mixed fractions into normal fractions by multiplying denominator by the non-fractional number and then adding the numerator. Put this quantity over the denominator.

OpenStudy (anonymous):

For example: \[6\frac{ 6 }{ 7 }=\frac{ 7*6+6 }{ 7 }=\frac{ 48 }{ 7 }\]

OpenStudy (anonymous):

Repeat this process for the rest